The average mpg usage for a 2004 Ford Expedition 2WD for a sample of 10 tanks of gas was 17.0
with a standard deviation of 0.8. For a Ford Explorer 2WD, the average mpg usage for a sample of
10 tanks of gas was 18.5 with a standard deviation of 1.0. (a) Assuming equal variances, at
α = .01, is the true mean mpg lower for the Ford Expedition? (b) Calculate the p-value using
Excel.

Answer by stanbon(75887)

The average mpg usage for a 2004 Ford Expedition 2WD for a sample of 10 tanks of gas was 17.0 with a standard deviation of 0.8.
---

For a Ford Explorer 2WD, the average mpg usage for a sample of
10 tanks of gas was 18.5 with a standard deviation of 1.0.


(a) Assuming equal variances, at α = .01, is the true mean mpg lower for the Ford Expedition?
Ho: u(Explor)-u(Exped) = 0
Ha: u(Explor)-u(Exped) > 0
---
I ran a 2-Sample T test and got
test statistic: t = 3.704
p-value = 0.0008685
df = 20-2 = 18
-----------------------
Conclusion Reject Ho. ; Exped < Explo
==========================================
Cheers,
stan H.
======================================